OpenStudy (anonymous):

how is .368 written as a fraction 8th grade

pooja195 (pooja195):

= 368/1000 = 46/125 Work: Write the decimal number as a fraction (over 1) 0.368 = 0.368 / 1 Multiplying by 1 to eliminate 3 decimal places we multiply top and bottom by 3 10's Numerator (N) N = 0.368 × 10 × 10 × 10 = 368 Denominator (D) D = 1 × 10 × 10 × 10 = 1000 N / D = 368 / 1000 Simplifying our fraction = 368/1000 = 46/125

OpenStudy (anonymous):

ty :) i guess im rlly dumb h